a new two contract for the building was signed with a 10.00% increase from the prior year
insurance expenses will hold steady due to contract agreement
travel/entertainment will be reduced by 7.00% from prior year
new equipment was purchased in the current year causing a 22.00% increase in depreciation
30% of the payroll cost will increase/decrease based upon the trend; 70% of the payroll will increase by 25.00%  
due to anticipated new hires causing contract labor to decrease by 30.00%
all the new hires will cause training expenses to increase by 11.00%
2014 2015 2016
Sales Revenue 9,800,000 10,780,000 12,397,000
Cost of Goods Sold 6,018,000     6,318,900     6,950,790
Gross Margin 3,782,000     4,461,100     5,446,210
Operating Expenses
Payroll Expenses 1,336,000     1,362,720     1,417,229
Building Expenses      284,800        284,800        284,800
Office Expenses      355,200        337,440        303,696
Advertising Expenses      844,800     1,056,000     1,320,000
Insurance Expenses      299,200        299,200        299,200
Contract Labor Expenses      149,600        127,160           95,370
Travel/Entertainment Expenses        89,500           90,395           91,299
Training Expenses        44,000           41,800           39,710
Service Charges Expenses        13,000           13,000           13,650
Dues/Licenses Expenses          9,800             9,800             9,800
Depreciation Expenses        52,450           49,828           47,336
Total Operating Expenses 3,478,350     3,672,143     3,922,090
Net Income      303,650        788,958     1,524,120
Based upon the assumptions listed above and using trend analysis for the  
remainder of the accounts, create a budget for 2017.
Answer #1
Budgeted Income statement
For the year 2017
Sales Revenue 13636700
Cost of goods sold 7298330
Gross Margin 6338370
Operating Expenses:
Payroll Expense 1690753
Building Expenses 284800
Office Expenses 349250
Advertising Expenses 1650000
Insurance Expenses 299200
Contract Labor Expenses 66759
Travel/Entertainment Expenses 84908
Training Expenses 44078
Service Charges Expenses 13650
Dues/Licenses Expenses 9800
Depreciation Expenses 57750
Total Operating Expenses 4550948
Net Income 1787422

Calculations:

Sales Revenue =12,397,000+10% 13636700
Insurance Expenses 299200
Travel/Entertainment Expenses = 91299-7% 84908
Depreciation = 47336+22% 57750
Payroll cost:
Trend [(1362720-1336000)/1336000]x100 = 2% (1417229-1362720)/1362720 = 4%
For 30% (1417229)x30%+6% 450678
For 70% (1417229x70%)+25% 1240075
Total 1690753
Contract labor = 95,370 - 30%(95,370) 66759
Training Expense = 39,710 + 11% 44078
2014 2015 2016 2017
Sales 9800000 10780000 12397000 13636700
Difference 980000 1617000 1239700
Trend 10% 15% 10%
Cost of goods sold 6018000 6318900 6950790 7298330
Difference 300900 631890 347540
Trend 5% 10% 5%
Office Expenses 355200 337440 303696 349250
Difference 17760 33744 45554
Trend 5% 10% 15%